OpenStudy (anonymous):

Find the area of a rectangle with length of 6 inches and width of 2 feet. In sq. ft.: In sq. in.:

OpenStudy (anonymous):

6in x (2x12in) = 144 sq in 6/12 ft x 2 ft = 1 sq ft

OpenStudy (anonymous):

2ft=24in A=6*24 A=144in 6in = 1/2ft A=(1/2)*2 A= 1ft